# Alternating Direction Method of Multipliers AR AutoRally

### v1

@@ 5             ${\bf x}\leftarrow{\bf x}_{0}$;
7                   ${\bf v}_{t}\leftarrow{\bf u_{t}}+\epsilon^{m}(t),\ \epsilon^{m}(t)\sim\mathcal{N}\left(0,\sigma^{2}\right)$;
@@ 9                   ${\bf y}\leftarrow G\left({\bf x},{\bf v}_{t}\right)$;
10                   $J^{m}\mathrel{+}=\ell\left({\bf y},{\bf v}_{t}\right)$;
@@ 13       $\eta\leftarrow\sum_{m=1}^{M}\exp\left(-\frac{1}{\lambda}\left(J^{m}-\rho\right)\right)$;
15             $w_{m}\leftarrow\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\left(J^{m}-\rho\right)\right)$;
@@
19             $\mathcal{U}_{t}\leftarrow{\bf u}_{t}+\sum_{m=1}^{M}w_{m}\epsilon^{m}(t)$;
@@
The original derivation of MPPI was done from a path-integral approach [4]. Future papers then derived MPPI from information theory [8], stochastic search [43], and mirror descent [44] approaches. A Tube-based MPPI controller [9] was also created in order to improve robustness to state disturbances. It made use of a tracking controller to cause the real system to track back to a nominal system that ignored state disturbances resulting in large costs. The nominal system would use the dynamics equation from Eq. 1 to update the initial state for a new iteration of MPPI instead of the real system’s state estimate. Both the real and the nominal trajectories are calculated using MPPI while the tracking controller was iLQR. In this setup, the tracking controller would always be the one sending controls to the system and as MPPI was not aware of the tracking controller, it could end up fighting against the tracking controller. In order to address this, RMPPI was developed in [45, 10], which applied the tracking controller feedback within the samples MPPI used. RMPPI also chose the initial state for the nominal system using a constrained optimization problem that tried to keep the nominal state as close to the real state without causing the resulting trajectory to have a cost larger than a given cost threshold $\alpha$. Choosing the nominal state in this way ensured there was an upper bound on how quickly the optimal trajectory’s cost could grow due to disturbance. Our library contains implementations of these algorithmic improvements to MPPI as different controllers are the best choice in different scenarios.
II-B Library Description
@@
The Dynamics and Cost Function classes are self-evident and are classes describing the ${\bf F}$, ${\bf G}$, $\ell$, and $\phi$ functions from Eqs. 1, 2 and 3. The Controller class finds the optimal control sequence $U$ that minimizes the cost ${\bf J}(Y,U)$ using algorithms such as MPPI. The Sampling Distributions are used by the Controller class to generate the control samples used for determining the optimal control sequence. The Feedback Controller class determines what feedback controller is used to help push the system back towards the desired trajectory. This is required for the Tube-MPPI and RMPPI implementations but can even be used with the MPPI controller as it provides feedback in between MPC iterations. Unless otherwise specified, the Feedback Controllers in code examples are instantiated but turned off by default. Finally, Plants are \@iaciMPC MPC wrapper around a given controller and are where the interface methods in and out of the controller are generally defined. For example, a common-use case of MPPI is on a robotics platform running Robot Operating System (ROS) [46]. The Plant would then be where you would implement your ROS subscribers to information such as state, ROS publishers of the control output, and the necessary methods to convert from ROS messages to MPPI-Generic equivalents. Each class type has their own parameter structures which encapsulate the adjustable parameters of each specific instantiation of the class. For example, the cartpole dynamics parameters include the mass and length of the pendulum whereas a double integrator dynamics system has no additional parameters.
@@ III Performance Implementation
We now will go over some of the performance-specific implementation details we make use of in MPPI-Generic. First, we will give a brief introduction to GPU hardware and terminology followed by general GPU performance tricks. We will then dive into computational optimizations that are specific to the tasks our library is used to solve.
@@ III-A GPU Parallelization Overview
The GPU is a highly parallelizable hardware device that performs operations a bit differently from a CPU. The lowest level of computation in CUDA is a warp, which consists of $32$ threads doing the same operation at every clock step. These warps are grouped together to produce thread blocks. The individual warps in the thread block are not guaranteed to be at the same place in the code at any given time and sometimes it can actually be more efficient to allow them to differ. The threads in a thread block all have access a form of a local cache called shared memory. Like any memory shared between multiple threads on the CPU, proper mutual exclusion needs to be implemented to avoid race conditions. Threads in a block can be given indices in 3 axes, $x$, $y$, and $z$, which we use to denote different types of parallelization within the library. Thread blocks can themselves be grouped into grids and are also organized into $x$, $y$, and $z$ axes. This is useful for large parallel operations that cannot fit within a single thread block and also do not need to use shared memory between every thread. The GPU code is compiled into kernels, which can then be provided arbitrary grid and block dimensions at runtime.
It is important to briefly understand the hierarchy of memory before discussing how to improve GPU performance. At the highest level, we have global memory which is generally measured in $\mathrm{GB}$s and very slow to access data from. Next, we have an L2 cache in the size of $\mathrm{MB}$s which can speed up access to frequently-used global data. Then we have the L1 cache, shared memory, and CUDA texture caches. The L1 cache and shared memory are actually the same memory on hardware and are generally several $\mathrm{kB}$s in size; they are just separated by programmers explicitly using shared memory and the GPU automatically filling the L1 cache. The CUDA texture cache is a fast read-only memory used for CUDA textures which are 2D or 3D representations of data such as a map.
@@ III-B General GPU speedups
When looking into writing more performant code, there are some general tricks that we have leveraged throughout our code library. The first is the use of CUDA streams [47]. By default, every call to the GPU blocks the CPU code from moving ahead. CUDA streams allow for the asynchronous scheduling of tasks on the GPU while the CPU performs other work in the meantime. We use CUDA streams throughout in order to schedule memory transfers between the CPU and GPU as well as kernel calls and have different streams for controller optimal control computation and visualization.
The next big tip is minimizing global memory accesses. Global memory reading or writing can be a large bottleneck in computation time and for our library, it can take up more time than the actual computations we want to do on the GPU. The first step recommended is to move everything you can from global memory to shared memory [48]. We also make use of Curiously Recurring Template Patterns [49] as our choice of polymorphism on the GPU as to avoid the need of constructing and reading from a virtual method table which would need to be stored in global memory. In addition, we utilize vectorized memory [50] accessing where possible. Looking at the GPU instruction set, CUDA provides instructions to read and write in 32, 64, and 128 bit chunks in a single instruction. This means that it is possible to load up to four $32$-bit floats in a single instruction. Using these concepts, we greatly reduce the number of calls to global memory and consequently increase the speed at which our computations can run.
Figure 1: Diagram of the execution flow of computeControl(). The blue ellipses indicate variables, the green rectangles are GPU methods, and the orange rectangles are CPU methods. The selection in purple is a single GPU kernel when using the combined kernel and separated out when using split kernels. Most of the code is run on the GPU but we have found that some operations such as finding the baseline run faster on the CPU.
@@ III-C Library-Specific Performance Optimizations
So far, we have discussed optimizations that can be done for any CUDA program. However, there are further optimizations to be had in choosing how to parallelize specific components of our library. In Fig. 1, we have the general steps taken every time we want to compute a new optimal control sequence in MPPI. These same steps are also taken in Tube-MPPI and RMPPI though they have to be done for both the nominal and real systems.
One major performance consideration is how to parallelize the Dynamics and Cost Function calculations. We have found that, depending on the circumstances and the number of samples used in MPPI, different parallelization techniques make more sense. One way would be to run the Dynamics and Cost Function in a combined kernel on the GPU while another would be to run them in separate kernels. We discuss the description as well as the pros and cons of each parallelization technique below.
III-C1 Split Kernel Description
We start by taking the initial state and control samples and run them through the Dynamics kernel. This kernel uses all three axes of thread parallelization for different components. First, the $x$ dimension of the block and the grid are used to indicate which sample we are on as threadIdx.x + blockDim.x * blockIdx.x. As every sample should go through the exact same computations, using the $x$ axis allows us to ensure that each warp is aligned. Next, the $z$ axis is used to indicate which system is being run; for MPPI, there is only one system but Tube-MPPI and RMPPI use two systems, nominal and real. Finally, the $y$ dimension is used to parallelize within the dynamics function. As dynamics are rarely doing the same derivative computation for every state, this additional parallelization within the dynamics, shown in Lst. 1, can lead to better performance rather than sequentially going through each calculation.
@@ III-C1 Split Kernel Description
In the Dynamics kernel, we then run a for loop over time for each sample in which we get the current control sample, runs it through the Dynamic’s step() method, and save out the resulting output to global memory.
@@
Now, we could have also run the Cost Function inside the previous kernel but we instead separate it out into its own kernel. The reason for that is that while the Dynamics must be sequential over time, the cost function does not need to be. To achieve this, we move the sample index up to the grid level and use the block’s $x$ axes for time instead. The Cost kernel gets the control and output corresponding to the current time in its computeRunningCost() or terminalCost() methods, adds the cost up across time for each sample, and saves out the resulting overall cost for each sample. A problem that might arise with this implementation is that we might become limited in the number of timesteps we could optimize over due there being a limit of $1024$ threads in a single thread block. In order to address this, we calculate the max number of iterations over the thread $x$ dimension required to achieve the desired number of timesteps and conduct a for loop over that iteration count. So for example, if we had $500$ as the desired number of timesteps and block $x$ size of $128$, we would do four iterations in our for loop to get the total horizon cost. These choices brings the time to do the cost calculation to much closer to that of a single timestep instead of having to wait for sequential iterations of the cost if it was paired with the Dynamics kernel.
III-C2 Combined Kernel Description
@@
The Combined Kernel runs the Dynamics and Cost Function methods together in a single kernel. This works by getting the initial state and control samples, applying the Dynamics’ step() to get the next state and output, and running that output through the Cost Functions’ computeCost() to get the cost at that time. This basic interaction is then done in a for loop over the time horizon to get the the entire state trajectory as well as the cost of the entire sample. We parallelize this over three axes. First, the $x$ and $z$ dimensions of the block and grid are used to indicate which sample and system we are on as described above in the Split Kernel’s Dynamics section. Finally, the $y$ dimension is used to parallelize within the Dynamics and Cost Functions’ methods.
III-C3 Choosing between the Split and Combined Kernels
@@
There are some trade-offs between the two kernel options that can affect the overall computation time. By combining the Dynamics and Cost Function calculations together, we can keep the intermediate outputs in shared memory and don’t need to save them out to global memory. However, we are forced to run the Cost Function sequentially in time. Splitting the Dynamics and Cost Function into separate kernels allows them each to use more shared memory for their internal calculations with the requirement of global memory usage to save out the sampled output trajectories. The Combined Kernel uses less global memory but requires more shared memory usage in a single kernel as it must contain both the Dynamics and Cost Functions’ shared memory requiremens. As the number of samples grow, the number of reads and writes of outputs to global memory also grows. This can eventually take more time to do than the savings we get from running the Cost Function in parallel across time, even when using vectorized memory reads and writes.
@@
In order to address these trade-offs, we have implemented both kernel approaches in our library and automatically choose the most appropriate kernel at Controller construction. The automatic kernel selection is done by running both the combined and split kernels multiple times and then choosing the fastest option. As the combined kernel potentially uses more shared memory than the split kernel, we also check to see if the amount of shared memory used is below the GPU’s shared memory hardware limit; if it is not, we default to the split kernel approach. We also allow the user to overwrite the automatic kernel selection through the use of the setKernelChoice() method.
III-C4 Weight Transform and Update Rule Kernels
@@
Once the costs of each sample trajectory is calculated, we then bring these costs back to the CPU in order to find the baseline, $\rho$. The baseline is calculated by finding the minimum cost of all the sample trajectories; it is subtracted out during the exponentiation stage as it has empirically led to better optimization performance. When the number of samples is only in the thousands, we have found that the copy to the CPU to do the baseline search is faster than attempting to do the search on the GPU.

### v2

@@ 4             ${\bf x}\leftarrow{\bf x}_{0}$;
6                   ${\bf v}_{t}\leftarrow{\bf u_{t}}+\epsilon^{m}_{t},\ \epsilon^{m}_{t}\sim\mathcal{N}\left(0,\Sigma\right)$;
@@ 8                   ${\bf y}\leftarrow G\left({\bf x},{\bf v}_{t}\right)$;
9                   $J^{m}\mathrel{+}=\ell\left({\bf y},{\bf v}_{t}\right)-\lambda\mathcal{I}\left({\bf v}_{t}\right)$;
@@ 12       $\eta\leftarrow\sum_{m=1}^{M}\exp\left(-\frac{1}{\lambda}\left(J^{m}-\rho\right)\right)$;
14             $w^{m}\leftarrow\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\left(J^{m}-\rho\right)\right)$;
15
@@ 14             $w^{m}\leftarrow\frac{1}{\eta}\exp\left(-\frac{1}{\lambda}\left(J^{m}-\rho\right)\right)$;
18             $\mathcal{U}_{t}\leftarrow{\bf u}_{t}+\sum_{m=1}^{M}w^{m}\epsilon^{m}_{t}$;
19
@@ 18             $\mathcal{U}_{t}\leftarrow{\bf u}_{t}+\sum_{m=1}^{M}w^{m}\epsilon^{m}_{t}$;
II-D Library Description
@@
The Dynamics and Cost Function classes are self-evident and are classes describing the ${\bf F}$, ${\bf G}$, $\ell$, and $\phi$ functions from Eqs. 1, 19 and 20. The Controller class finds the optimal control sequence ${\bf U}^{*}$ that minimizes the cost in Eq. 20 using algorithms such as MPPI. The Sampling Distributions are used by the Controller class to generate the control samples used for determining the optimal control sequence. The Feedback Controller class determines what feedback controller if any, is used to help push the system back towards the desired trajectory computed by the Controller. This is required for the Tube-MPPI and RMPPI implementations but can even be used with the MPPI controller as it provides feedback in between MPC iterations. Unless otherwise specified, the Feedback Controllers in code examples are instantiated but turned off by default. Finally, Plants are \@iaciMPC MPC wrapper around a given controller and are where the interface methods in and out of the controller are generally defined. For example, a common-use case of MPPI is on a robotics platform running Robot Operating System (ROS) [47]. The Plant is where you would implement your ROS subscribers to information such as state, ROS publishers of the control output, and the necessary methods to convert from ROS messages to MPPI-Generic equivalents. Each class type has their own parameter structures which encapsulate the adjustable parameters of each specific instantiation of the class.
@@ III Performance Implementation
We shall now discuss some of the performance-specific implementation details we make use of in MPPI-Generic. First, we will give a brief introduction to GPU hardware and terminology followed by general GPU performance tricks useful for implementation in this library.
@@ III-A GPU Parallelization Overview
The GPU is a highly parallelizable hardware device that performs operations differently than a CPU. The lowest level of computation in CUDA is a warp, which consists of $32$ threads doing the same operation at every clock step [48]. These warps are grouped together to produce thread blocks. While the threads in a warp are computed together, the individual warps in the thread block are not guaranteed to be at the same place in the code at any given time and sometimes it can actually be more efficient to allow them to differ. The threads in a thread block all have access a form of a local cache called shared memory. Like any memory shared between multiple threads on the CPU, proper mutual exclusion needs to be implemented to avoid race conditions. Threads in a block can be given indices in 3 axes, $x$, $y$, and $z$, which we use to denote different types of parallelization within the library. The conversion from a thread’s 3D index of $(x,y,z)$ to its thread number in the block is given by (z * blockDim.y + y) * blockDim.x + x. Thread blocks can themselves be grouped into grids and are also organized into $x$, $y$, and $z$ axes. This is useful for large parallel operations that cannot fit within a single thread block.The GPU code is compiled into kernels, which can be provided arbitrary grid and block dimensions at runtime.
It is important to briefly understand the hierarchy of memory before discussing how to improve GPU performance. At the highest level, we have global memory which is generally measured in $\mathrm{GB}$s and very slow to access data from. Next, we have an L2 cache in the size of $\mathrm{MB}$s which can speed up access to frequently-used global data. Then we have the L1 cache, shared memory, and CUDA texture caches. The L1 cache and shared memory are actually the same memory on hardware and are generally several $\mathrm{kB}$s in size; they are separated by programmers explicitly using shared memory and the GPU automatically filling the L1 cache. The CUDA texture cache is a fast read-only memory used for CUDA textures which are 2D or 3D representations of data such as a map.
@@ III-B General GPU speedups
When looking into writing more performant code, there are some general tricks that we leveraged throughout our code library. The first is the use of CUDA streams [49]. By default, every call to the GPU blocks the CPU code from moving ahead. CUDA streams allow for the asynchronous scheduling of tasks on the GPU while the CPU performs other work in the meantime. We use CUDA streams throughout in order to schedule memory transfers between the CPU and GPU as well as kernel calls and have different streams for controller optimal control and visualization computations.
The next big tip is minimizing global memory accesses. Global memory reading or writing can be a large bottleneck in computation time and for our library, it can be slower than the actual computations we want to do on the GPU. The first recommendation is to move commonly-accessed data from global memory to shared memory [50]. We also use Curiously Recurring Template Patterns [51] as our choice of polymorphism on the GPU to avoid the need of constructing and reading from a virtual method table which would be stored in global memory.
We utilize vectorized-memory [52] accessing where possible. Looking at the GPU instruction set, CUDA provides instructions to read and write in 32, 64, and 128 bit chunks, making it possible to load up to four $32$-bit floats in a single instruction. Using these concepts, we greatly reduce the number of calls to global memory and consequently increase the speed at which our computations can run.
We also make use of hardware-defined mathematical operators called intrinsics in some places as well to reduce computation time [53]. These instrinsics are approximations of various mathematical operations such as division, $\sin$, etc. that are implemented at the hardware level, requiring many fewer clock cycles to compute. The trade-off is that they are approximations and can return incorrect evaluations depending on the inputs. As such, we limited our use of intrinsics to trigonometric functions like __cosf(), __sinf(), and __sincosf(). We found that other intrinsics such as __fdividef() or __expf() when used throughout the code base cause significantly different optimal control sequence calculations. However, there are plenty of specific locations in the code where more intrinsics can be introduced without negative effects on accuracy at a future point. In addition, we try to make use of float-specific methods when applicable such as expf() to prevent unnecessary conversions to and from double.
Figure 1: Diagram of the execution flow of computeControl(). The blue ellipses indicate variables, the green rectangles are GPU methods, and the orange rectangles are CPU methods. The selection in purple is a single GPU kernel when using the combined kernel and separated out when using split kernels. Most of the code is run on the GPU but we found that some operations such as finding the baseline and calculating the normalizer, $\eta$, run faster on the CPU.
@@ III-C Library-Specific Performance Optimizations
So far, we discussed optimizations that can be done for any CUDA program. However, there are further optimizations to be had in choosing how to parallelize specific components of our library. In Fig. 1, we have the general steps taken every time we want to compute a new optimal control sequence in MPPI. These same steps are also taken in Tube-MPPI and RMPPI though they have to be done for both the nominal and real systems.
One major performance consideration is how to parallelize the Dynamics and Cost Function calculations. We can either run the Dynamics and Cost Function in a combined kernel or run them in separate kernels. We describe each parallelization technique as well as the pros and cons below. First, we introduce a slight modification to the Dynamics and Cost Functions.
III-C1 Intermediate Calculation Passthrough
When creating Cost Functions for a given Dynamics, it might be required to redo calculations already done in the Dynamics. For example, putting a penalty on the location of the wheels of a vehicle inherently requires knowing where the wheels are located. The wheel locations can be calculated given the position and orientation of the center of mass of the vehicle and so are not considered part of the state. Depending on the Dynamics, the wheel locations might also be calculated as part of the state update. To reduce unnecessary recalculations, we provide a way to pass these extra values directly from the Dynamics to the Cost Function. We do this by slightly modifying Eq. 2 to use outputs, ${\bf y}_{t}$ instead of ${\bf x}_{t}$,
${\bf y_{t}}$ $={\bf G}\left({\bf x}_{t},{\bf u}_{t}\right)$ (19)
where ${\bf G}\left({\bf x}_{t},{\bf u}_{t}\right)$ is calculated as part of the Dynamics. For the vast majority of systems, ${\bf y}_{t}$ is the true state, i.e. ${\bf y}_{t}={\bf x}_{t}$, but we have found in some cases that bringing additional calculations from the Dynamics to the Cost Function can be computationally faster than reproducing them.
III-C2 Split Kernel Description
We start by taking the initial state and control samples and run them through the Dynamics kernel. This kernel uses all three axes of thread parallelization for different components. First, the $x$ dimension of the block and the grid are used to indicate which sample we are on as threadIdx.x + blockDim.x * blockIdx.x. As every sample will conduct the exact same computations, using the $x$ axis allows us to ensure that each warp is aligned as long as the $x$ block size is chosen appropriately. Next, the $z$ axis is used to indicate which system is being run; for MPPI, there is only one system but Tube-MPPI and RMPPI use two systems, nominal and real. As dynamics generally have different derivative computations for different states, we use the $y$ dimension, as shown in Lst. 1, to introduce additional parallelization within the dynamics, instead of sequential computation of each state derivative, which can lead to further performance improvements. When our thread block’s $x$ dimension is a multiple of 32, the $y$ threads are separated into different warps and Lst. 1 improves performance. However, when a switch/if statement causes threads in the same warp to follow different computations, this is known as warp divergence, and the GPU runs the warp again to go through all code paths. Depending on the complexity of the branching, this can cause significant slowdowns.
@@ III-C2 Split Kernel Description
In the Dynamics kernel, we run a for loop over time for each sample in which we get the current control sample, run it through the Dynamic’s step() method, and save out the resulting output to global memory.
@@
Next, we look to the Cost Function ran inside its own kernel. The reason for that is that while the Dynamics must be sequential over time, the cost function does not need to be. To achieve parallelization across time, we move the sample index up to the grid level and use the block’s $x$ axes for time instead. The Cost kernel gets the control and output corresponding to the current time in its computeRunningCost() or terminalCost() methods, adds the cost up across time for each sample, and saves out the resulting overall cost for each sample. A problem that arises is the limited number of timesteps we could optimize over due to the limit of $1024$ threads in a single thread block; we address this by reusing threads to each compute multiple timesteps until we reach the desired time horizon. These choices brings the time to calculate the cost much closer to that of a single timestep instead of having to wait for sequential iterations of the cost if it was paired with the Dynamics kernel.
III-C3 Combined Kernel Description
@@
The Combined Kernel runs the Dynamics and Cost Function methods together in a single kernel. This works by getting the initial state and control samples, applying the Dynamics’ step() to get the next state and output, and running that output through the Cost Functions’ computeCost() to get the cost at that time. This basic interaction is done in a for loop over the time horizon to get the the entire state trajectory as well as the cost of the entire sample. We parallelize this over three axes. First, the $x$ and $z$ dimensions of the block and grid are used to indicate which sample and system we are on as described above in the Split Kernel’s Dynamics section. Finally, the $y$ dimension is used to parallelize within the Dynamics and Cost Functions’ methods.
III-C4 Choosing between the Split and Combined Kernels
@@
There are some trade-offs between the two kernel options that can affect the overall computation time. By combining the Dynamics and Cost Function calculations together, we keep the intermediate outputs in shared memory and do not need to save them out to global memory. However, we are forced to run the Cost Function sequentially in time. Splitting the Dynamics and Cost Function into separate kernels allows them each to use more shared memory for their internal calculations with the requirement of global memory usage to save out the sampled output trajectories. The Combined Kernel uses less global memory but requires more shared memory usage in a single kernel as it must contain both the Dynamics and Cost Functions’ shared memory requirements. As the number of samples grow, the number of reads and writes of outputs to global memory also grows. This can eventually take more time than the savings we get from running the Cost Function in parallel across time, even when using vectorized-memory reads and writes.
@@
In order to address these trade-offs, we implemented both kernel approaches in our library and automatically choose the most appropriate kernel in the Controller constructor using chooseAppropriateKernel(). The automatic kernel selection is done by running both the combined and split kernels multiple times, measuring the computation time of each, and choosing the fastest option. As the combined kernel potentially uses more shared memory than the split kernel, we also check to see if the amount of shared memory used is below the GPU’s shared memory hardware limit; if it is not, we default to the split kernel approach. We also allow the user to overwrite the automatic kernel selection through the use of the setKernelChoice() method.
III-C5 Weight Transform and Update Rule Kernels
@@
Once the costs of each sample trajectory is calculated, we bring these costs back to the CPU in order to find the baseline, $\rho$. The baseline is calculated by finding the minimum cost of all the sample trajectories; it is subtracted out during the exponentiation stage as it has empirically led to better optimization performance. When the number of samples is only in the thousands, we found that the copy to the CPU to do the baseline search is faster than doing so on the GPU. The costs are then exponentiated on the GPU and the normalizer, $\eta$, is calculated back on the CPU before doing the final optimal control calculation.